the square has side length of 2x+9 the perimeter is 68 whats the value of x
Answer by jim_thompson5910(35256)

Perimeter = 4*(Side Length)

P = 4*s

P = 4*(2x+9)

68 = 4*(2x+9)

4*(2x+9) = 68

2x+9 = 68/4

2x+9 = 17

2x = 17-9

2x = 8

x = 8/2

x = 4

The value of x is x = 4